Problem 25.25

Part A

What is the electric field amplitude of an electromagnetic wave whose magnetic field amplitude is 2.9mT ?
Answer #1

In Harmonic electromagnetc wave a direct relationship between electric and magnetic field exists.Electromagnetic Wave Propagation

E=Emcos(kx-wt)

B=Bmcos(kx-wt)
integrating Electric field with x and magnetic field with time yields
dE/dx=-kEmsin(kx-wt)

dB/dt=wBmsin(kx-wt)

Maxwells 3rd equation leads to
dE/dx=-dB/dt
=>kEm=wBm or Bm=Em/c
Magnetic field amplitude is 2.9mT
So, Em=Bm*c=2.9*10-3*3*108=870000v/m or 87*104v/m